can someone help me with this question?
If f(x)=x^2+3x-5 then f(-2)=

this was a question developed in class. thanks

Answer by checkley75(3666)

IF F(X)=X^2+3X-5 THEN IF F(-2)=
F(-2)=-2^2+3(-2)-5
F(-2)=4-6-5
F(-2)=-7